OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (mathstudent55):

Add equation 1 to equation 2. Since you have -y in eq 1 and y in eq 2, the y's will cancel out and you can solve for x easily. x - y = 16 2x + y = 2 --------- 3x = 18 Divide both sides by 3 x = 6 Now substitute x into eq 1: 6 - y = 16 Subtract 6 from both sides: -y = 10 divide both sides by -1 y = -10

OpenStudy (radar):

I would choose substitution method as the first equation is simple to use: x=y+16 substituting this in second equation it then becomes 2(y+16) + y = 2 2y +32 +y = 2 3y +32 = 2 3y = -30 y = -10

OpenStudy (mathstudent55):

The above was part 1. Part 2 is Elimination method because by adding the equations as they are you eliminate y and it's easy to solve for x.

OpenStudy (radar):

Then substitute 10 for y in the first equation getting: x-(-10) =16 x+10=16 x=6

OpenStudy (anonymous):

Thank you guys for your help! :)